OpenStudy (anonymous):

How do I convert between logarithmic and exponential form? I know how to do it, but how would I explain it in a general form?

OpenStudy (zehanz):

You need what is in fact the definition of the logarithm: \(\log_{a}b=c \Leftrightarrow a^c=b\)

OpenStudy (mertsj):

\[\log_{b}y=x means b^x=y\]

OpenStudy (zehanz):

@Mertsj: if you want to get a space in Latex, just type a tilde: ~

OpenStudy (anonymous):

how do you know?

OpenStudy (anonymous):

does that apply to all probs?

OpenStudy (anonymous):

@ZeHanz

OpenStudy (zehanz):

I know, because I had to learn it, just like you :) It is what a logarithm is: it is an exponent. I.e. \(\log_2 8=3\) because \(2^3=8\).

OpenStudy (anonymous):

wow...so intelligent

OpenStudy (zehanz):

 *blushing*